OpenStudy (anonymous):

Verify that quadrilateral BCDE is a rhombus with vertices B(5, -1), C(-2, 0), D( -1, 7) and E(6, 6) by showing that all four sides are equal.

OpenStudy (anonymous):

These are my answer choices A. BD = CD = \[\sqrt{50}\] B. BD = CE = 50 C.BC = CD = DE = EB = 50 D. BC = CD = DE = EB = \[\sqrt{50}\]

OpenStudy (anonymous):

How would you solve this? Please help!

jimthompson5910 (jim_thompson5910):

To find the length of each side, you use the distance formula This works because the distance from B to C is the same as the length of segment BC

jimthompson5910 (jim_thompson5910):

Distance formula \[\LARGE d = \sqrt{\left(x_1 - x_2\right)^2+\left(y_1 -y_2\right)^2}\]

jimthompson5910 (jim_thompson5910):

For example B = (x1,y1) = (5,-1) so x1 = 5 and y1 = -1 C = (x2,y2) = (-2,0) so x2 = -2 and y2 = 0 \[\LARGE d = \sqrt{\left(x_1 - x_2\right)^2+\left(y_1 -y_2\right)^2}\] \[\LARGE d = \sqrt{\left(5 - (-2)\right)^2+\left(-1-0\right)^2}\] \[\LARGE d = \sqrt{\left(5 +2\right)^2+\left(-1-0\right)^2}\] \[\LARGE d = \sqrt{\left(7\right)^2+\left(-1\right)^2}\] \[\LARGE d = \sqrt{49+1}\] \[\LARGE d = \sqrt{50}\] \[\LARGE d = \sqrt{25*2}\] \[\LARGE d = \sqrt{25}*\sqrt{2}\] \[\LARGE d = 5\sqrt{2}\] So the exact distance from B to C is \(\Large 5\sqrt{2}\) units The approximate length of BC (ie the approx distance) is 5*sqrt(2) = 7.0710678

jimthompson5910 (jim_thompson5910):

do the same for the other side lengths

OpenStudy (anonymous):

Thank you! @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you're welcome